Recent Log Entries

Hall B current ramp rate vs current spikes

Log Number: 4066185Logbook(s): Entry Makers: kpriceBody: 

A recent spate of detector trips in Hall B may possibly be traced back to current spiking upon beam restoration. In an effort to try and reduce the peak current, I turned the ramp rate down from 1.5 nA/sec to 0.8 nA. Not sure if this reduction was the thing that helped, but Hall reports fewer detector trips for the time being.

open-loop tuned 1L14-5

Log Number: 4066155Logbook(s): Entry Makers: kpriceBody: 

Cavity still has a ragged GMES. Looking through the archiver it seems this has always been the case, regardless of GSET. Tried open-loop tuning anyways and found that it was close to resonance but there might be something nonstandard going on with its tuner response.

GMES is still not very steady; taking more gradient out of it probably isn't a great idea unless we re-lem.

iCalibrate: Hall C

Log Number: 4066144Logbook(s): Body: Calibration ParametersHall:CTarget: Loop 1 10cmPass: 5Note: DeuteriumCalibrated On:2022-Sep-18 13:41Calibrated By: lrichardFile:C-2022-09-18_1341_Loop_1_10cm_Pass-5.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Dump Left</t

Follow-up Re: 2L21-3 Stuck SOSing - High KBCU

Log Number: 4066139Logbook(s): Body: 

This was noted Yesterday afternoon (10/21/22). Issue is an instrumentation / readback problem as it has not affected beam operation. The screen shows faults on the RF Quick Page for that RFCM. Normal operating parameters are shown for the cavity. Problem to be addressed at the first opportunity.

Hall A -> 45 uA on polarized He3

Log Number: 4066104Logbook(s): Entry Makers: kpriceBody: 

Hall A requests 45 uA CW on polarized He3 glass cell. Raster is on (amplitude 1.3409 V, 1.2068 H). Ramp rate protection is on. Ion chambers are set for 45 uA with a 25% margin. MaxJuice is set to 50 uA.

Beam authorization: up to 60 uA CW
Operational restrictions for glass cells: up to 60 uA CW with raster always on

iCalibrate: Hall A

Log Number: 4066099Logbook(s): Body: Calibration ParametersHall:ATarget: Reference Cell H2 100pctPass: 2Note: 6mm RasterCalibrated On:2022-Oct-16 21:54Calibrated By: bschaumlFile:A-2022-10-16_2154_Reference_Cell_H2_100pct_Pass-2.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)

iCalibrate: Hall C

Log Number: 4066087Logbook(s): Body: Calibration ParametersHall:CTarget: Loop 2 10cmPass: 5Note: Most recentCalibrated On:2022-Oct-01 13:22Calibrated By: lrichardFile:C-2022-10-01_1322_Loop_2_10cm_Pass-5.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Dump Left<

Operator SIGN IN

Log Number: 4066057Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Price
DATE:             10/22/2022
SHIFT:               Day

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

Operator SIGN IN

Log Number: 4066056Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Blackshaw
DATE:             10/22/2022
SHIFT:               Day

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

Total Beam Power

Log Number: 4066041Logbook(s): Entry Makers: chumphryBody: 

HA: 45 uA HB: 8 nA HC: 40 uA HD: 200 nA

Autolog save

Log Number: 4066040Logbook(s): Tags: Body: saver ran from opsl632

Saved: chumphry 10-22-22 07:43:56 {A:2 B:5 C:5 E:1047 A=CW B=CW C=CW D=CW 0R08=82.14 No dumps in.

Pages